# Active particle motion in Poiseuille flow through rectangular channels

### v1

@@ Acknowledgments
Figure 12: Comparison of (a) simpler flow field approximation $\bar{u}(x,y)$ as in Eq. (8) used in this paper with (b) a more accurate flow field $\bar{u}_{e}(x,y)$ as in Eq. (17) with $n=100$ terms for Poiseuille flow in a $3$D square duct. The difference $\bar{u}_{e}(x,y)-u(x,y)$ is shown in panel (c).
Figure 13: Classification of active particle trajectories in the initial condition $(x(0),y(0))$ plane using a more accurate flow field in the square cross-section $\bar{u}_{e}(x,y)$ (as per Eq. (17) with $n=100$ terms) compared to the simpler flow field $\bar{u}(x,y)$ used in the paper as in Fig. 4. We note that the grid resolution here is $50\times 50$ compared to the $100\times 100$ resolution in Fig. 4. Other parameters were fixed to $\theta(0)=0$, $\phi(0)=0$ and $\bar{U}=10$.
@@ Appendix A Calculation for general constant of motion
For dynamics of the active particle suspended in a unidirectional flow field $\bar{u}(x,y)$ as given by Eqs. (3), we show that following is a constant of motion:
$H_{g}=-\frac{1}{2}\bar{u}(x,y)+e_{z}.$
@@
$\frac{\text{d}H_{g}}{\text{d}t}=-\frac{1}{2}\frac{\partial\bar{u}}{\partial x}-\frac{1}{2}\frac{\partial\bar{u}}{\partial y}+0+0+0+\left(\frac{1}{2}\frac{\partial\bar{u}}{\partial x}+\frac{1}{2}\frac{\partial\bar{u}}{\partial y}\right)=0.$
@@ Appendix B Comparison of exact versus approximate flow field for Poiseuille flow in a square cross-section
The dimensionless flow field $\bar{u}(x,y)$ used in this paper is an approximation to the following exact flow field for Poiseuille flow in a straight duct with a square cross-section [32]:
$\bar{u}_{e}(x,y)=\bar{U}_{e}(1-y^{2})$ (17)
$+\frac{32\bar{U}_{e}}{\pi^{3}}\sum_{n=0}^{\infty}\frac{(-1)^{n+1}\cosh\left({(2n+1)\frac{\pi x}{2}}\right)\,\cos\left({(2n+1)\frac{\pi y}{2}}\right)}{(2n+1)^{3}\,\cosh\left({(2n+1)\frac{\pi}{2}}\right)}.$
Here $\bar{U}_{e}=\bar{U}/\max\{\bar{u}_{e}(x,y)\}=\bar{U}/{\bar{u}_{e}(0,0)}$, to match the maximum flow speed of $\bar{U}$ at the center of the channel. A comparison of the flow field $\bar{u}$ and $\bar{u}_{e}$ for $\bar{U}=1$ is shown in Fig. 12. We see that the overall qualitative flow field is captured well by our approximate flow field $\bar{u}$ and the difference between the two flow fields is small; at most $6\%$ when scaled by the maximum flow speed at the center. Further, we note that the regions in the cross-section where the most significant difference is observed in the flow field are near the corners of the square.
Figure 13 shows the different types of active particle trajectories realized for a more accurate flow field $\bar{u}_{e}(x,y)$ for Poiseuille flow in a straight channel with square cross-section. Comparing with Fig. 4(a), which used a simpler approximation of the flow field $\bar{u}(x,y)$, we find noticeable differences mainly in the yellow regions. This is to be expected since at these locations, the difference between the flow fields $\bar{u}_{e}$ and $\bar{u}$ is the largest (see Fig. 12(c)).

### v2

@@ Acknowledgments
Figure 12: Comparison of (a) the flow field ${u}(x,y)$ as in Eq. (7) used in this paper to approximate Poiseuille flow in a $3$D rectangular channel (here setting $AR=1$ for a square channel) with (b) the more accurate representation of the flow field ${u}_{e}(x,y)$ as in Eq. (15) with $n=100$ terms. The difference ${u}_{e}(x,y)-u(x,y)$ is shown in panel (c).
Figure 13: Classification of active particle trajectories in the initial condition $(x(0),y(0))$ plane using the more accurate flow field in the square cross-section ${u}_{e}(x,y)$ (as per Eq. (15) with $n=100$ terms). This compares well with the classification shown in Fig. 4, obtained using the approximate flow field ${u}(x,y)$ with $AR=1$. We note that the grid resolution here is $50\times 50$ compared to the $100\times 100$ resolution in Fig. 4. Other parameters were fixed to $\theta(0)=0$, $\phi(0)=0$ and $U=10$.
@@ Appendix A Calculation for general constant of motion
For dynamics of the active particle suspended in a unidirectional flow field ${u}(x,y)$ as given by Eqs. (2), we show that following is a constant of motion:
$H_{g}=-\frac{1}{2}{u}(x,y)+e_{z}.$
@@
$\frac{\text{d}H_{g}}{\text{d}t}=-$ $\frac{1}{2}\frac{\partial{u}}{\partial x}e_{x}-\frac{1}{2}\frac{\partial{u}}{\partial y}e_{y}+0+0+0$
$+$ $\left(\frac{1}{2}\frac{\partial{u}}{\partial x}e_{x}+\frac{1}{2}\frac{\partial{u}}{\partial y}e_{y}\right)=0.$
@@ Appendix B Comparison of exact versus approximate flow field for Poiseuille flow in a square cross-section
The dimensionless flow field ${u}(x,y)$ used in this paper is an approximation to the following exact flow field for Poiseuille flow in a straight duct with a square cross-section [32]:
${u}_{e}(x,y)={U}_{e}(1-y^{2})$ (15)
$+\frac{32{U}_{e}}{\pi^{3}}\sum_{n=0}^{\infty}\frac{(-1)^{n+1}\cosh\left({(2n+1)\frac{\pi x}{2}}\right)\,\cos\left({(2n+1)\frac{\pi y}{2}}\right)}{(2n+1)^{3}\,\cosh\left({(2n+1)\frac{\pi}{2}}\right)}.$
Here ${U}_{e}={U}/\max\{{u}_{e}(x,y)\}={U}/{{u}_{e}(0,0)}$, to match the maximum flow speed of ${U}$ at the center of the channel. A comparison of the flow field ${u}$ and ${u}_{e}$ for ${U}=1$ is shown in Fig. 12. We see that the overall qualitative flow field is captured well by our approximate flow field ${u}$ and the difference between the two flow fields is small; at most $6\%$ when scaled by the maximum flow speed at the center. Further, we note that the regions in the cross-section where the most significant difference is observed in the flow field are near the corners of the square. For wider rectangular cross-sections, the approximation becomes poorer with increasing AR. For a $2\times 1$ rectangular cross-section the maximum difference between the two velocity fields is around $15\%$ while for a $4\times 1$ rectangular cross-section it is around $35\%$.
Figure 13 shows the different types of active particle trajectories realized for the more accurate flow field ${u}_{e}(x,y)$ for Poiseuille flow in a straight channel with square cross-section. Comparing with Fig. 4(a), which used the simpler approximation of the flow field ${u}(x,y)$, we find noticeable differences mainly in the yellow regions. This is to be expected since at these locations, the difference between the flow fields ${u}_{e}$ and ${u}$ is the largest (see Fig. 12(c)).